OpenStudy (ehall):

Which option equals tan 40°? Select all that apply. sin50/cos50 cos50/sin50 cos40/sin40 sin40/cos40

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

tangent is sine over cosine, so one answer is D

OpenStudy (misty1212):

there is another answer, since \(\sin(50)=\cos(40)\) and vice versa

OpenStudy (ehall):

thank you for helping me

OpenStudy (misty1212):

don't forget your other answer \[\frac{\cos(50)}{\sin(50)}\]

OpenStudy (ehall):

thank you so much because I am new to this open study

OpenStudy (misty1212):

\[\color\magenta\heartsuit\]